OpenStudy (clamin):

HELP!! MEDAL!! A biology class takes a field trip to a nearby...field, and collects data on the biodiversity (the number of different plant and animals species) in their local ecosystem. Each student randomly selects one square meter of ground, and counts all the different species of plants, animals bugs, mold, etc that are within that square meter, as well as the total numbet of organisms. Question 1: Student A finds 27 unique species in her square meter, and decides to use Fermi Estimation to speed up the rest of her data collection. If she counts 23 grass plants in 100 square centimeter

OpenStudy (clamin):

s, about how many grasses are there in her entire square meter? [Assume the grass is evenly distiributed]

OpenStudy (jtvatsim):

We can use unit factor label conversions to solve \[\frac{23 \text{ grass plants}}{100 \text{ cm}^2} \times \frac{100 \text{ cm}}{\text{m}} \times \frac{100 \text{ cm}}{\text{m}} = \frac{230,000 \text{ grass plants}}{\text{m}^2}\]